An exclusive yoghurt manufacturer sells 4,000 gallons per month at a price of 40 cedis each. When the price is reduced to 30 cedis sales increase to 6,000 gallons per month

a, calculate the price elasticity of demand for the yogurts over this price range


b, Is demand elastic unit elastic or inelastic


c, Calculate the change in revenue due to the change in price


1
Expert's answer
2021-05-04T11:17:03-0400

Percentage change in quantity "\Delta Q=\dfrac{6000-4000}{4000}\times 100=50" %


Percentage change in price "\Delta P=\dfrac{40-30}{40}\times 100=25" %


(i)  Price elasticity of demand "=\dfrac{\Delta Q}{\Delta P}=\dfrac{50}{25}=2"


(ii) As The price elasticity of demand is greater than 1, So Demand is elastic.


(iii)Change in Total revenue "=(6000\times 30)-(4000\times 40)=180000-160000=20000" ced.